OpenStudy (anonymous):

Rewrite each equation in slope-intercept form. Graph each line. 22. 4x-2y=3

OpenStudy (anonymous):

4x - 2y = 3 -4x -4x -2y = -4x +3 /-2 /-2 y = 2x -3/2 

OpenStudy (dayakar):

3=4x-2y 3+2y=4x 2y= 4x-3 y=( 4x-3)/2

OpenStudy (dayakar):

y= 4x/2 -3/2 y= 2x -3/2 y= 2x -1.5

OpenStudy (dayakar):

it is in the form of y= mx+b y=2x -1.5 compare both slope (m) =2 y-intercept (b) = -1.5

OpenStudy (dayakar):



OpenStudy (dayakar):

can u draw a graph with above two points